# MFTF: Mask-free Training-free Object Level Layout Control Diffusion Model

### v1

@@ Abstract
Text-to-image generation models have become transformative tools. However, diffusion-based vision language models still lack the ability to precisely control the shape, appearance, and positional placement of objects in generated images using text guidance alone. Global image editing models typically achieve global layout control by relying on additional masks or images as guidance, which often require model training. While local object editing models enable modification of object shapes, they do not provide control over the positional placement of these objects. To address these limitations, we propose the MFTF model, which enables precise control over object positioning without requiring additional masks or images. The MFTF model supports both single-object and multi-object positional control (such as translation, rotation, etc.) and allows for concurrent layout control and object semantic editing. This is achieved by controlling the denoising process of the diffusion model through parallel denoising. Attention masks are dynamically generated from the cross-attention layers of the source diffusion model and applied to queries from the self-attention layers to isolate objects. These queries are then modified according to layout control parameters and injected back into the self-attention layers of the target diffusion model to enable precise positional control.
Figure 1: Our model successfully achieves both single-object and multi-object layout control, as well as simultaneous layout control and semantic editing, without image guidance or model training and fine-tuning. The images are generated using the source prompt $P_{s}$, with additional layout control parameters $L$ applied, alongside concurrent semantic editing via the target prompt.
@@ 1 Introduction
Text-to-image generation models have emerged as transformative tools, enabling the production of high-quality, photorealistic, and artistically sophisticated visuals from natural language inputs. Advances in vision-language generative modeling primarily fall into two categories: transformer-based autoregressive approaches [3, 25] and diffusion-based methods [18, 7, 19, 11]. Among these, diffusion models have shown outstanding promise, delivering state-of-the-art results across various benchmarks [10, 15, 17, 14, 1].
@@
Diffusion based vision language models, however, lack the capability to precisely control the generated images using text guidance alone [1]. Although images generated from the same text prompt with different initial noise maintain semantic consistency, they often exhibit considerable variations in shape, appearance, and object positioning [12].
@@
Current layout control methods encompass global layout control and local object editing. Global layout control typically utilizes additional masks or images as guidance and may require model training, depending on whether auxiliary networks are employed [20, 26]. While existing local object editing techniques provide control over semantics and structure, they still lack the ability to achieve precise control over the positional placement of objects [6, 12, 2, 8].
@@
To complement existing spatial control techniques, we propose MFTF model, which enables precise positional control of individual objects without requiring additional masks as guidance or model training, shown in Fig. 1. The model achieves this by controlling the denoising process of the diffusion model through parallel denoising for both the source and target prompts. Attention masks are dynamically generated from the cross-attention layers of the source diffusion model and applied to queries from the self-attention layers to isolate objects. These queries are then modified according to layout control parameters and injected back into the self-attention layers of the target diffusion model, enabling precise positional control throughout the denoising process. Additionally, the cross-attention masks generated for tokens can also be used for text-guided semantic segmentation. The architecture of the proposed model is illustrated in Fig. 2.
The challenge lies in performing object-level layout control without relying on additional image guidance, and object background seperation. Drawing inspiration from previous work [6, 2, 20] that modifies either the cross-attention layer or the self-attention layer, as well as employing multi-self-attention techniques, our model generates attention masks dynamically. These masks are used in the cross-attention layers to effectively separate objects from the background within the self-attention layers. Ultimately, layout control of the objects is achieved through a modified multi-self-attention operation, allowing for precise control over the positioning and arrangement of the objects in the generated images.
The challenge lies in performing object-level layout control without relying on additional image guidance and object-background separation. Inspired by previous work [6, 2, 20] that modifies either the cross-attention layer or the self-attention layer, and utilizes multi-self-attention techniques, our model dynamically generates attention masks, applied in the self-attention layers to effectively separate objects from the background. Ultimately, object layout control is achieved through a modified multi-self-attention operation.
@@
The contributions of this model are as follows: (1) The MFTF model enables single-object and multi-object level positional control, as well as concurrent object-level layout control and semantic editing. (2) MFTF supports concurrent layout control and semantic editing without requiring images or masks as guidance without model training or fine-tuning. (3) The proposed method can also be applied to text-controlled semantic segmentation.
@@ 2.1 Text-to-Image Generation
Text-to-image generation primarily consists of two families of models: auto-regressive models and diffusion models. Among the auto-regressive models, the Parti [24] model integrates both text and image modules using a transformer architecture, enabling text-controlled, high-fidelity, photorealistic image generation. The Parti model is built upon VQVAE [21] and VQGAN [5, 23], which facilitate the tokenization and detokenization of images, which enables the fusion of text tokens and image tokens within the transformer architecture [22].
The diffusion-based family of text-to-image generation models has gained significant attention in recent years. Notable models in this category include Imagen [17], which utilizes pretrained large language models for text encoding, achieving strong text-image alignment and photorealistic image generation. DALL-E 2,3 [14, 1] from OpenAI achieve text and image alignment through the CLIP [13] model and generate images using a diffusion model. The Latent Diffusion Model (LDM) adopts a different architecture from DALL-E 2,3 by conditioning on text for image generation, achieving text-image alignment within the latent space [15]. Our model builds upon the open-source Stable Diffusion model, which is an open-source implementation of the Latent Diffusion Model.
@@ 2.2 Text-to-Image Generation Layout Control and Semantic Editing
While local object editing models enable shape modification of objects, they lack the ability to control the positional placement of objects. The Prompt-to-Prompt method [6] enables object editing while preserving the original object positions by modifying the text prompt. This method identifies the corresponding cross-attention layers for text tokens and swaps, inserts, or re-weights the attention maps for those tokens. Patashnik et al. [12] applied prompt mixing combined with image-space localization techniques to edit the shape of local objects while maintaining the layout, semantics, and textures of the images. Masactrl [2] achieved object editing by modifying cross-attention and self-attention layers but did not alter the positioning of objects. Imagic [8] performed local object editing through text embedding optimization and diffusion model fine-tuning.
Global semantic editing and texture editing models that achieve layout control in diffusion models often rely on additional images or masks as guidance, with some requiring model training and others not. The Plug-and-Play model [20] facilitates semantic editing while preserving the layout of the guidance image, which is achieved by injecting features into the self-attention layers. ControlNet [26], on the other hand, enhances layout control by incorporating ”zero convolutions” and using external inputs such as instance edges, depth maps, segmentation, or human poses to manipulate the global image layout. However, ControlNet necessitates model training for its application.
To address these limitations, our MFTF model enables precise control over the position of objects without requiring additional masks or images as guidance. In addition to supporting single-object and multi-object positional control (including translation, rotation, etc.), MFTF model facilitates concurrent layout control and object semantic editing. MFTF model complements existing methods by enabling more flexible positional control in image generation. Furthermore, it can be applied for text-guided semantic segmentation.
@@ 3.1 Latent Diffusion Model
The diffusion model is inspired by principles of non-equilibrium thermodynamics [18, 7, 19, 11] and consists of two complementary processes: the forward process and the reverse process. In the forward process $q(x_{t}|x_{t-1})$, Gaussian noise $\epsilon$ is incrementally added to the original image $x_{0}$, progressively transforming it into random noise $x_{T}\sim N(0,I)$. In contrast, the reverse process $p_{\theta}(x_{t-1}|x_{t})$ systematically removes the noise $\epsilon$ from the noisy image $x_{t}$, ultimately reconstructing the original image.
Our model is based on the latent diffusion model [15], which enhances computational efficiency by performing diffusion in the latent space. This approach reduces the time complexity of both the forward and reverse processes while preserving synthesis quality [15]. Latent diffusion model employs either a variational autoencoder (VAE) [9, 4] or a VQGAN [5] for encoding and decoding images. Specifically, the encoder $\varepsilon$ maps the input image $x$ to a latent representation $z$, while the decoder $D$ reconstructs $z$ back to the image domain. The diffusion processes, relying on a U-Net architecture [16], are executed entirely in the latent space $z$. To incorporate conditional guidance, a spatial transformer is integrated into the U-Net, enabling enhanced flexibility. Additionally, conditions are encoded using the encoder $\tau_{\theta}$ and incorporated into the diffusion model via cross-attention layers within the spatial transformer architecture. This design ensures robust and effective guidance throughout the diffusion process.
@@ 3.2 Cross-attention and Self-attention in Diffusion Models
In the stable diffusion model, the modified U-Net module $\epsilon_{\theta}$ incorporates a spatial transformer block, which consists of a self-attention layer followed by a cross-attention layer. During the denoising process, the encoded image $z$ is combined with the time-step embedding $t$ and passed through the self-attention layer. The output from the self-attention layer is then processed by the cross-attention layer, where it interacts with text-guidance embeddings. The formulations of self-attention and cross-attention are presented in Eq. 1 and Eq. 2, respectively. Here, $Q_{z},K_{z}$ and $V_{z}$ are derived from spatial features, $K_{\tau},V_{\tau}$ are derived from text embeddings, and $d$ represents the embedding dimension.
@@
The self-attention and cross-attention layers encode rich semantic and structural information, which can be leveraged to control style and semantic concepts, as demonstrated in Prompt-to-Prompt [6] and Plug-and-Play [20]. Additionally, the cross-attention mechanism can be utilized to generate masks for separating objects from the background, as implemented in Masactrl [2]. Specifically, objects from token $i$ can be isolated from the background by applying the cross-attention mask $M^{l}_{i}$. This enables layout control through the application of the layout parameter $L$, which governs structural transformations such as rotation, position translation, and scaling, thereby facilitating precise control over the generated layout.
@@ 4 Simultaneous Object Level Layout Control and Semantic Editing
The primary objective of this model is to facilitate simultaneous object-level layout control and semantic editing without the need for additional image or mask guidance, and without requiring model training or fine-tuning. As demonstrated in Fig. 1, the model effectively achieves positional control of individual objects, enabling layout transformations such as translation, scaling, and rotation to be applied concurrently. Moreover, it illustrates that both object-level layout control and semantic editing can be performed simultaneously. Additionally, the model supports the control of multiple object positions concurrently. To realize this objective, we propose the MFTF model, which generates attention masks and utilizes cross-attention mechanisms to separate objects from the background within the self-attention layers. Object layout control is then achieved through multi-self-attention [2].
@@
The architecture of the model is illustrated in Fig. 2. The model utilizes a prompt $P_{s}$ for generating the source image $x_{s}$, and a separate prompt $P_{t}$ for generating the target image $x_{t}$. The target prompt $P_{t}$ can either remain the same as $P_{s}$ or differ, depending on whether semantic editing is applied to the image. The object indices $P^{i}_{s}$ from $P_{s}$ and the object layout control parameters $L^{i}$ are provided to the model to regulate the layout of objects in the generated target image $x_{t}$. For example, as shown in Fig. 1, the images are generated with for instance the position of the ”castle” is shifted to the left according to $L$, the “steam” in $P_{s}$ is replaced with “desert” in $P_{t}$.
Dive into details of the MFTF model. It governs the denoising process of the diffusion model by facilitating parallel denoising for both the source and target prompts. It dynamically generates attention masks from the cross-attention layers of the source diffusion model. These attention masks are then applied to the queries $Q_{s}$ derived from the self-attention layers to identify the objects. Subsequently, the modified queries $Q_{ms}$ are generated in accordance with the layout control parameters. Finally, $Q_{ms}$ is injected into the self-attention layer of the target diffusion model, enabling precise positional control during the denoising process.
Figure 2: MFTF Architecture. MFTF dynamically generates attention masks from the cross-attention layers of the source diffusion model. These attention masks are then applied to the queries $Q_{s}$ derived from the self-attention layers to separate the objects from background. Subsequently, the modified queries $Q_{ms}$ are generated in accordance with the layout control parameters. Finally, $Q_{ms}$ is injected into the self-attention layer of the target diffusion model, enabling precise positional control during the denoising process.
@@
The MFTF algorithm is summarized in Algorithm 1. The parameter $t^{*}$ represents the cutoff time step for applying layout control, while $T$ denotes the total number of denoising time steps. For time steps $t<t^{*}$, the mask $M^{i}_{s}$ is obtained from the cross-attention layer, with $Q_{s}$ derived from the self-attention layer. Details of the cross-mask generation are provided in Sec. 4.1. The mask $M^{i}_{s}$ is then applied to the query $Q_{s}$ in the self-attention mechanism to separate the object from the background. The object’s position is adjusted according to the given layout control parameters $L$. Subsequently, the modified query $Q_{ms}$ is passed to the target diffusion model to generate the target latent space $Z_{t}$. Details of the self-attention query $Q_{ms}$ generation are further explained in Sec. 4.1.
@@
The MFTF model enables simultaneous object-level layout control and semantic Editing and through three key mechanisms: (1) the cross-attention mask, which encodes semantic and structural information of objects for token/object $i$, despite potential overlaps in semantic features; (2) the query $Q$ from the self-attention layers, which incorporates layout information and structural details of the generated images, allowing positional control by modifying the objects within $Q_{s}$; and (3) the diffusion model generate structural details during the later stages of the denoising process.
@@
4.1 Dynamic Cross-attention Mask Generation
The cross-attention mask is constructed utilizing the cross-attention layer, where the object mask is derived following the formula described in Eq. 3, with a specified threshold $eta$. The generated masks, containing both the object and the background, are depicted in Fig. 3.
The cross-attention map, defined as $A^{i}_{s}=QK_{T}$, is represented by the dimensions $b,n,m$, where $b$ corresponds to the batch size, $n$ denotes the feature sequence length, and $m$ represents the text sequence length. The model is capable of dynamically generating cross-attention masks $M^{i}_{s}$ for token $i$ at each sampling step $t$, which are applied to the self-attention layer $l$. This mechanism allows for controlled image generation using dynamically updated cross-attention masks.

### v2

@@ Abstract
Text-to-image generation models have revolutionized content creation, but diffusion-based vision-language models still face challenges in precisely controlling the shape, appearance, and positional placement of objects in generated images using text guidance alone. Existing global image editing models rely on additional masks or images as guidance to achieve layout control, often requiring retraining of the model. While local object-editing models allow modifications to object shapes, they lack the capability to control object positions. To address these limitations, we propose the Mask-free Training-free Object-Level Layout Control Diffusion Model (MFTF), which provides precise control over object positions without requiring additional masks or images. The MFTF model supports both single-object and multi-object positional adjustments, such as translation and rotation, while enabling simultaneous layout control and object semantic editing. The MFTF model employs a parallel denoising process for both the source and target diffusion models. During this process, attention masks are dynamically generated from the cross-attention layers of the source diffusion model and applied to queries from the self-attention layers to isolate objects. These queries, generated in the source diffusion model, are then adjusted according to the layout control parameters and re-injected into the self-attention layers of the target diffusion model. This approach ensures accurate and precise positional control of objects. Project source code available at https://github.com/syang-genai/MFTF.
Figure 1: MFTF successfully achieves single-object and multi-object layout control, as well as simultaneous layout control and semantic editing, without guidance image or mask and no model training and fine-tuning.
@@ 1 Introduction
Text-to-image generation models have become powerful tools for creating high-quality, photorealistic, and artistically refined images from natural language descriptions. Recent advances in vision-language generative modeling can be broadly categorized into two approaches: transformer-based autoregressive methods [3, 25] and diffusion-based techniques [18, 7, 19, 11]. Among these, diffusion models have demonstrated exceptional performance, achieving state-of-the-art results across a wide range of benchmarks [10, 15, 17, 14, 1].
@@
However, diffusion-based vision-language models struggle to provide precise control over the generated images using text guidance alone [1]. While images generated from the same text prompt with different initial noise typically maintain semantic consistency, they often display significant variations in object shape, appearance, and positioning [12].
@@
Current layout control methods can be broadly categorized into global layout control and local object editing. Global layout control typically relies on additional masks or images as guidance, and may require model retraining depending on whether auxiliary networks are used [20, 26]. While existing local object editing techniques allow for control over semantics and structure, they still fall short of enabling precise control over the positional placement of objects [6, 12, 2, 8].
@@
To enhance existing spatial control techniques, we introduce MFTF model, which allows for precise positional control of individual objects without the need for additional masks or model retraining, as illustrated in Fig. 1. The architecture of the proposed model is shown in Fig. 2. The model achieves this by manipulating the denoising process of the diffusion model through parallel denoising for both the source and target prompts. Attention masks are dynamically generated from the cross-attention layers of the source diffusion model and applied to queries from the self-attention layers to isolate objects. These modified queries are then adjusted based on layout control parameters and reintroduced into the self-attention layers of the target diffusion model, enabling accurate positional control throughout the denoising process. Furthermore, the cross-attention masks generated for tokens can be leveraged for text-guided semantic segmentation in SFig. 11.
@@
The challenges in achieving object-level layout control lie in doing so without relying on additional image guidance. Building on previous work [6, 2, 20] that modifies either the cross-attention or self-attention layers and employs multi-self-attention techniques, our model dynamically generates attention masks. These masks are applied within the self-attention layers to effectively distinguish objects from the background. Ultimately, object layout control is accomplished through a modified multi-self-attention operation in supplement Fig. 8. The contributions of this model are as follows: (1) The MFTF model enables single-object and multi-object level positional control, as well as concurrent object-level layout control and semantic editing. (2) MFTF supports concurrent layout control and semantic editing without requiring images or masks as guidance without model training or fine-tuning. (3) The proposed method can also be applied to text-controlled semantic segmentation.
Figure 2: MFTF Architecture. MFTF dynamically generates attention masks from the cross-attention layers of the source diffusion model. These attention masks are then applied to the queries $Q_{s}$ derived from the self-attention layers to separate the objects from background. Subsequently, the modified queries $Q_{ms}$ are generated in accordance with the layout control parameters $L$. Finally, $Q_{ms}$ is injected into the self-attention layer of the target diffusion model, enabling precise positional control via denoising process.
@@ 2.1 Text-to-Image Generation
Text-to-image generation primarily encompasses two families of models: autoregressive models and diffusion models. Among the autoregressive models, the Parti [24] model integrates both text and image modules using a transformer architecture, enabling text-controlled, high-fidelity, and photorealistic image generation. Built upon VQVAE [21] and VQGAN [5, 23], the Parti model leverages these methods to tokenize and detokenize images, allowing for the fusion of text and image tokens within the transformer framework [22].
The diffusion-based family of text-to-image generation models has garnered significant attention in recent years. Notable models in this category include Imagen [17], which leverages pretrained large language models for text encoding, achieving strong text-image alignment and photorealistic image generation. OpenAI’s DALL-E 2 and 3 [14, 1] utilize the CLIP [13] model to align text and images, generating visuals via a diffusion model. The Latent Diffusion Model (LDM) employs a distinct architecture from DALL-E 2 and 3, conditioning on text for image generation and achieving text-image alignment within the latent space [15]. Our model builds upon the open-source Stable Diffusion model, which is an open-source implementation of the Latent Diffusion Model.
Figure 3: Cross-attention mask $M^{i}_{s}$ is generated from the cross-attention layers $A^{i}_{s}$
Figure 4: Visualization of $Q$ is presented at different self-attention layers $l=11,13,15$, showing masks from the following conditions: without cross-attention mask, with cross-attention mask applied, and with additional positional control $L$.
@@ 2.2 Text-to-Image Generation Layout Control and Semantic Editing
While local object editing models enable shape modification of objects, they lack the ability to control the positional placement of objects. The Prompt-to-Prompt method [6] enables object editing while preserving the original object positions. Patashnik et al. [12] applied prompt mixing combined with image-space localization techniques to edit the shape of local objects while maintaining the layout, semantics, and textures of the images. Masactrl [2] achieved object editing by modifying cross-attention and self-attention layers but still can not alter the positioning of objects. Imagic[8] performed local object editing through text embedding optimization and diffusion model fine-tuning.
Global semantic editing and texture editing models that achieve layout control in diffusion models often rely on additional images or masks as guidance, with some requiring model training and others not. The Plug-and-Play model [20] facilitates semantic editing while preserving the layout of the guidance image, which is achieved by injecting features into the self-attention layers. ControlNet [26], on the other hand, enhances layout control by incorporating ”zero convolutions” and using external inputs such as instance edges, depth maps, segmentation, or human poses to manipulate the global image layout. However, ControlNet necessitates model training.
To address these limitations, our MFTF model enables precise control over the position of objects without requiring additional masks or images as guidance and no training. In addition to supporting single-object and multi-object positional control, including translation, rotation, etc.. It facilitates concurrent layout control and object semantic editing. MFTF model complements existing methods by enabling more flexible positional control in image generation.
@@ 3.1 Latent Diffusion Model
Diffusion model is inspired by principles of non-equilibrium thermodynamics [18, 7, 19, 11] and consists of two complementary processes: the forward process and the reverse process. In the forward process $q(x_{t}|x_{t-1})$, Gaussian noise $\epsilon$ is incrementally added to the original image $x_{0}$, progressively transforming it into random noise $x_{T}\sim N(0,I)$. In contrast, the reverse process $p_{\theta}(x_{t-1}|x_{t})$ systematically removes the noise $\epsilon$ from the noisy image $x_{t}$, ultimately reconstructing the original image.
Our model builds upon the latent diffusion model [15], which improves computational efficiency by performing the diffusion process in the latent space $z$. This approach significantly reduces the time complexity for forward and reverse processes while maintaining high-quality image synthesis [15]. Then latent diffusion model utilizes either a variational autoencoder (VAE) [9, 4] or a VQGAN [5] for image encoding and decoding. Specifically, the encoder maps an input image $x$ into a compact latent representation $z$, and the decoder reconstructs $z$ back to the image domain $x$. The diffusion process, powered by a U-Net architecture [16], operates entirely in the latent space $z$. To enable conditional guidance, a spatial transformer is integrated into the U-Net, providing enhanced flexibility. Additional conditions are encoded using an encoder $\tau_{\theta}$ and incorporated into the diffusion model through cross-attention layers within the spatial transformer. This architecture ensures robust and effective guidance throughout the diffusion process, facilitating high-quality and controlled image generation.
@@ 3.2 Cross-attention and Self-attention in Diffusion Models
In Stable Diffusion model, the modified U-Net module $\epsilon_{\theta}$ integrates a spatial transformer block, which includes a self-attention layer followed by a cross-attention layer. The mathematical formulations for self-attention and cross-attention are provided in Eq. 1 and Eq. 2, respectively. In these formulations, $Q_{z}$, $K_{z}$, and $V_{z}$ are computed from spatial features in Eq. 1, $K_{\tau}$ and $V_{\tau}$ are derived from text embeddings in Eq. 2, $d$ denotes the embedding dimension.
@@
The self-attention and cross-attention layers capture rich semantic and structural information, enabling their use for controlling style and semantic concepts [6, 20]. Additionally, the cross-attention mechanism can generate masks to separate objects from the background, as employed in Masactrl [2]. MFTF applies the cross-attention mask $M^{l}_{i}$, objects corresponding to token $i$ can be effectively isolated from the background in self-attention layers.
Figure 5: Quantitative comparison of images: objective-level layout control via text descriptions vs. MFTF model; without vs. with MFTF model.
@@ 4 Simultaneous Object Level Layout Control and Semantic Editing
The primary aim of this model is to enable simultaneous object-level layout control and semantic editing without relying on additional image or mask guidance, as well as without necessitating model training or fine-tuning. Our MFTF model accomplishes this by dynamically generating attention masks and leveraging cross-attention mechanisms to separate objects from the background in the self-attention layers. Object layout control is subsequently realized through the application of modified multi-self-attention operations SFig. 8. As illustrated in Fig. 1, MFTF model achieves precise positional control of individual objects, allowing layout transformations such as translation, scaling, and rotation to be applied concurrently. Furthermore, the model demonstrates the capability to perform object-level layout control and semantic editing simultaneously while also supporting the concurrent manipulation of multiple object positions.
@@
The architecture of the proposed model is depicted in Fig. 2. The model operates using a source prompt $P_{s}$ to generate source image $x_{s}$ and a target prompt $P_{t}$ to generate target image $x_{t}$. The target prompt $P_{t}$ can either be identical to $P_{s}$ or differ, depending on whether semantic editing is applied. To control the layout of objects in the generated target image $x_{t}$, the model requires the object tokens $P^{i}_{s}$ from $P_{s}$ and the object layout control parameters $L^{i}$. For instance, as illustrated in Fig. 1, the model adjusts the position of the ”a crystal-clear stream” to the center based on $L$, and then replacing the ”forest” in $P_{s}$ with ”desert” in $P_{t}$.
@@
MFTF algorithm is summarized in Algorithm 1. The parameter $t^{*}$ represents the cutoff time step for applying layout control, while $T$ denotes the total number of denoising time steps. For time steps $t<t^{*}$, the mask $M^{i}_{s}$ is obtained from the cross-attention layer, with $Q_{s}$ derived from the self-attention layer. The cross-attention mask $M^{i}_{s}$ is then applied to the query $Q_{s}$ in the self-attention to separate the object from the background. Details of the cross-mask generation are provided in Sec. 4.1. The object’s position is adjusted according to the given layout control parameters $L$. Subsequently, the modified query $Q_{ms}$ is passed to the target diffusion model to generate the target latent space $z_{t}$. Details of the self-attention query $Q_{ms}$ generation are further explained in Sec. 4.2.
@@
MFTF model enables simultaneous object-level layout control and semantic editing and through three key mechanisms: (1) the cross-attention mask contains semantic and structural information of tokens, despite potential overlaps in semantic features (SFig. 9); (2) query $Q$ from the self-attention layers incorporates layout information and structural details of the generated images, allowing positional control by modifying the objects within $Q_{s}$ (3) the diffusion model generate structural details during the later stages of the denoising process.
@@
4.1 Cross-Attention Mask Creation
The cross-attention mask is constructed utilizing the cross-attention layer, where the object mask is derived following the formula described in Eq. 3, with a specified threshold $\eta$. The generated masks, containing both the object and the background, are depicted in Fig. 3. The threshold $\eta$ has a significant impact on the generated images through the attention mask. When $\eta=0$, no attention mask is applied to the query, making it impossible to separate the object from the background. When $\eta=1$, the query is entirely removed from the self-attention process. For multiple-object position control, the threshold $\eta$ controls the degree to which the original object is retained.